The goal of the shift is send the IR beam retroreflected by the filter cavity on EQB1.
After the recovery of the optical lever we realigned the filter cavity with the green beam:
- We misaligned by 1000 urad in TX the input mirror and we centered the direct beam in the FCEB_GR_CAM
- We religned the input mirror and we realigned the end mirror by 1000 urad in TX
- We superposed the incoming beam and the retroreflected beam on EQB1 looking after the green FI
- We realigned the end mirror and we sligthly moved the two filter cavity mirrors in order to maximize the flashes (max value found ~2V) see Fig 1.
When the cavity was aligned with the green beam we started to align the IR retroreflection up to EQB1
- Acting on Picomotors on SQB1 M21 and M22 we steered the IR retroreflection in order to have it transmitted by the FI
- When the retroreflection was transmitted by the FI we moved SQB1_M13 and SQB1_M14 in order to send the retroreflection on EQB1
- We have to move the diaphragm after SQB1_M14 because the IR beam was not centered into its hole.
After this operations we had the retroreflection on EQB1 and we started to align it into the homodyne cameras.
When we was doing the camera alignment we realized that the retroreflection was slightly clipped by the FI thus we moved M21 in order to unclip it. We measured the power of retroreflected IR beam after the FI and we had 1.208 mW compared by the 1.25mW of incoming beam.
After the realignment the retroreflected beam was too low on SQB1_M12 thus in the next shifts maybe could be useful to increase the height of the beam by acting both on M21 and M22
We tried to see IR flashes in transmission by the filter cavity without success. We didn't check the alignment of IR sensors and we didn't move the FLT Mirrors during this attempt.